# What’s Inside Your Diffusion Model? A Score-Based Riemannian Metric to Explore the Data Manifold

### v2

@@ B.1 Discretization and Numerical Integration
Given endpoints $\boldsymbol{x}_{A}$ and $\boldsymbol{x}_{B}$, we represent the geodesic path as a sequence of $n+1$ points:
@@
where $\gamma_{0}=\boldsymbol{x}_{A}$ and $\gamma_{n}=\boldsymbol{x}_{B}$ are fixed, and the interior points $\{\gamma_{1},\ldots,\gamma_{n-1}\}$ are optimized.
@@
    Compute the velocity vector: $\boldsymbol{v}_{i}=\gamma_{i+1}-\gamma_{i}$
@@
    Calculate the midpoint: $\boldsymbol{m}_{i}=\frac{1}{2}(\gamma_{i+1}+\gamma_{i})$
@@
    Evaluate the score function at the midpoint: $\boldsymbol{s}_{i}=\boldsymbol{s}(\boldsymbol{m}_{i})$
@@
$\mathcal{E}[\gamma]\approx\frac{1}{2}\sum_{i=0}^{n-1}\|\boldsymbol{v}_{i}\|^{2}_{g(\boldsymbol{m}_{i})}=\frac{1}{2}\sum_{i=0}^{n-1}\left[\|\boldsymbol{v}_{i}\|^{2}+\lambda(\boldsymbol{s}_{i}^{T}\boldsymbol{v}_{i})^{2}\right]$ (13)
@@
B.2 Riemannian Adam Optimization
@@
3:Initialize: Moment vectors $\boldsymbol{m}_{i}=\boldsymbol{0}$, $\boldsymbol{v}_{i}=\boldsymbol{0}$ for $i\in\{1,\ldots,n-1\}$
4:$t\leftarrow 0$
@@
8:     for $i=1$ to $n-1$ do
9:         $\boldsymbol{g}_{i}\leftarrow\nabla_{\gamma_{i}}\mathcal{E}_{\text{total}}$
10:         $\widetilde{\boldsymbol{g}}_{i}\leftarrow$ RiemannianGradient$(\boldsymbol{g}_{i},\gamma_{i},g)$triangleright Convert Euclidean to Riemannian gradient
11:         $\boldsymbol{m}_{i}\leftarrow\beta_{1}\boldsymbol{m}_{i}+(1-\beta_{1})\widetilde{\boldsymbol{g}}_{i}$
12:         $\boldsymbol{v}_{i}\leftarrow\beta_{2}\boldsymbol{v}_{i}+(1-\beta_{2})\widetilde{\boldsymbol{g}}_{i}^{2}$
13:         $\hat{\boldsymbol{m}}_{i}\leftarrow\boldsymbol{m}_{i}/(1-\beta_{1}^{t})$triangleright Bias correction
14:         $\hat{\boldsymbol{v}}_{i}\leftarrow\boldsymbol{v}_{i}/(1-\beta_{2}^{t})$
15:         $\boldsymbol{\eta}_{i}\leftarrow\alpha\cdot\hat{\boldsymbol{m}}_{i}/(\sqrt{\hat{\boldsymbol{v}}_{i}}+\epsilon)$triangleright Update direction
@@
17:         $\gamma_{i}\leftarrow\gamma_{i}-\boldsymbol{\eta}_{i}$triangleright Update position
18:         $\boldsymbol{m}_{i}\leftarrow$ ParallelTransport$(\boldsymbol{m}_{i},\gamma_{i}^{\text{old}},\gamma_{i},g)$triangleright Transport momentum vector
@@
Riemannian Gradient: The Euclidean gradient $\boldsymbol{g}$ is converted to a Riemannian gradient $\widetilde{\boldsymbol{g}}$ using:
$\widetilde{\boldsymbol{g}}=g(\boldsymbol{x})^{-1}\boldsymbol{g}$ (17)
For our metric tensor $g(\boldsymbol{x})=\mathbf{I}+\lambda\cdot\boldsymbol{s}(\boldsymbol{x})\boldsymbol{s}(\boldsymbol{x})^{T}$, we can efficiently compute this using the Sherman-Morrison formula:
$\widetilde{\boldsymbol{g}}=\boldsymbol{g}-\frac{\lambda\cdot(\boldsymbol{s}^{T}\boldsymbol{g})\cdot\boldsymbol{s}}{1+\lambda\cdot(\boldsymbol{s}^{T}\boldsymbol{s})}$ (18)
@@
[…]
@@
Appendix C Geodesic Computation Algorithm
@@
1:Input: Points $\boldsymbol{p}$, $\boldsymbol{q}$, diffusion timestep $t$, number of segments $n$, scale parameter $\lambda$
2:Output: Geodesic path $\gamma=\{\gamma_{0},\gamma_{1},\ldots,\gamma_{n}\}$ with $\gamma_{0}=\boldsymbol{p}$, $\gamma_{n}=\boldsymbol{q}$
@@
4:Sample noise $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$
5:$\tilde{\boldsymbol{p}}\leftarrow\sqrt{\alpha_{t}}\boldsymbol{p}+\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}$
6:$\tilde{\boldsymbol{q}}\leftarrow\sqrt{\alpha_{t}}\boldsymbol{q}+\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}$
@@
8:$\gamma_{i}\leftarrow(1-\lambda_{i})\tilde{\boldsymbol{p}}+\lambda_{i}\tilde{\boldsymbol{q}}$ for $i\in\{0,1,\ldots,n\}$, $\lambda_{i}=i/n$
@@
11:$g_{\boldsymbol{x}}(\boldsymbol{u},\boldsymbol{v})=\boldsymbol{u}^{T}\boldsymbol{v}+\lambda\cdot(\boldsymbol{s}(\boldsymbol{x})^{T}\boldsymbol{u})(\boldsymbol{s}(\boldsymbol{x})^{T}\boldsymbol{v})$
@@
17:     for $i=0$ to $n-1$ do
18:         $\boldsymbol{v}_{i}\leftarrow\gamma_{i+1}-\gamma_{i}$triangleright Segment velocity
19:         $\boldsymbol{m}_{i}\leftarrow\frac{1}{2}(\gamma_{i+1}+\gamma_{i})$triangleright Segment midpoint
20:         $\mathcal{E}\leftarrow\mathcal{E}+\frac{1}{2}\|\boldsymbol{v}_{i}\|^{2}+\frac{\lambda}{2}(\boldsymbol{s}(\boldsymbol{m}_{i})^{T}\boldsymbol{v}_{i})^{2}$triangleright Energy contribution
@@
23:     $\mathcal{E}_{\text{smooth}}\leftarrow\lambda_{\text{smooth}}\sum_{i=1}^{n-1}\|\gamma_{i+1}-2\gamma_{i}+\gamma_{i-1}\|^{2}$triangleright Smoothness
24:     $\mathcal{E}_{\text{mono}}\leftarrow\lambda_{\text{mono}}\sum_{i=1}^{n-1}\text{ReLU}(\|\gamma_{i}-\gamma_{n}\|-\|\gamma_{i-1}-\gamma_{n}\|)$triangleright Monotonicity
@@
35:         $\hat{\gamma}_{i}\leftarrow$ original clean point ($\boldsymbol{p}$ or $\boldsymbol{q}$)
@@
37:         $\hat{\gamma}_{i}\leftarrow$ Denoise($\gamma_{i}$, from $t$ to 00)
@@ Appendix D Interpolation
1:Input: Points $\boldsymbol{p}$, $\boldsymbol{q}$, number of interpolation points $n$, timestep $t$
2:Output: Interpolated sequence $\{\boldsymbol{y}_{0},\boldsymbol{y}_{1},\ldots,\boldsymbol{y}_{n}\}$
@@
4:Sample noise $\boldsymbol{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$triangleright Same noise for consistency
5:$\tilde{\boldsymbol{p}}\leftarrow\sqrt{\alpha_{t}}\boldsymbol{p}+\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}$
6:$\tilde{\boldsymbol{q}}\leftarrow\sqrt{\alpha_{t}}\boldsymbol{q}+\sqrt{1-\alpha_{t}}\boldsymbol{\epsilon}$
@@
16:         $\boldsymbol{y}_{i}\leftarrow\boldsymbol{p}$triangleright Use original clean endpoint
@@
18:         $\boldsymbol{y}_{i}\leftarrow\boldsymbol{q}$triangleright Use original clean endpoint
@@
20:         $\boldsymbol{y}_{i}\leftarrow$ DenoisingDiffusion($\tilde{\gamma}_{i}$, from $t$ to 00)
@@
23:return $\{\boldsymbol{y}_{0},\boldsymbol{y}_{1},\ldots,\boldsymbol{y}_{n}\}$
Appendix E Extrapolation
@@
1:Initialize $\boldsymbol{x}_{\text{current}}\leftarrow\boldsymbol{q}$
@@
6:     Compute score at current point: $\boldsymbol{s}\leftarrow\mathbf{s}(\boldsymbol{x}_{\text{current}})$
7:     Compute direction: $\boldsymbol{d}\leftarrow(1-\varepsilon)\cdot\boldsymbol{m}+\varepsilon\cdot\boldsymbol{s}$
[…]
9:     Update position: $\boldsymbol{x}_{\text{next}}\leftarrow\boldsymbol{x}_{\text{current}}+\boldsymbol{d}$
10:     Update momentum: $\boldsymbol{m}\leftarrow\beta\cdot\boldsymbol{m}+(1-\beta)\cdot\boldsymbol{d}$
11:     $\boldsymbol{x}_{\text{current}}\leftarrow\boldsymbol{x}_{\text{next}}$
12:     Add $\boldsymbol{x}_{\text{current}}$ to extrapolation path
@@
Appendix F Rotated MNIST
After setting up the dataset as in Section 4.2, we trained a DDPM [10] Model in PyTorch (version 2.7 and cuda 12.6) by exploiting the Diffusers library [39] and Accellerate library [40] to parallelize training on 7 NVIDIA A100 GPUs.
@@ Appendix F Rotated MNIST
  • Input/Output: The model accepts 32$\times$32 grayscale images (single channel) and predicts noise with matching dimensions.
@@
Appendix G Stable Diffusion & MorphBench
We employed the pre-trained Stable Diffusion 2.1. implementation available in the Diffusers library [39], running with PyTorch 2.7 and cuda 12.6 on a NVIDIA A100 GPU (40GB VRAM).
@@ Appendix G Stable Diffusion & MorphBench
For our experiments with Stable Diffusion 2.1, we used unconditional generation by providing an empty text prompt (""), allowing the model to focus solely on the image manifold structure without text-based guidance.
@@
Additionally we report more examples of interpolation comparison, see Figure 5,  6,  7
@@
avg_ssim $\uparrow$ 0.6251 $\pm$ 0.1084 $0$$\pm$ 0.1198 0.4202 $\pm$ 0.1152 0.6261 $\pm$ 0.1026
avg_psnr $\uparrow$ 21.23 $\pm$ 2.08 20.72 $\pm$ 2.37 17.09 $\pm$ 2.31 20.98 $\pm$ 2.05
avg_lpips $\downarrow$ 0.3467 $\pm$ 0.0749 0.3652 $\pm$ 0.0816 0.4577 $\pm$ 0.0584 0.3402 $\pm$ 0.0679
fid $\downarrow$ 125.89 $\pm$ 44.13 142.08 $\pm$ 49.26 253.52 $\pm$ 54.07 127.29 $\pm$ 49.95
kid $\downarrow$ 0.0947 $\pm$ 0.0652 0.1059 $\pm$ 0.0716 0.1925 $\pm$ 0.0791 0.0901 $\pm$ 0.0659
@@
avg_ssim $\uparrow$ 0.6282 $\pm$ 0.1290 0.5653 $\pm$ 0.1517 0.3969 $\pm$ 0.1201 0.6148 $\pm$ 0.1222
avg_psnr $\uparrow$ 21.29 $\pm$ 1.69 20.21 $\pm$ 2.19 16.58 $\pm$ 1.89 20.83 $\pm$ 1.65
avg_lpips $\downarrow$ 0.3672 $\pm$ 0.0774 0.4076 $\pm$ 0.0779 0.4891 $\pm$ 0.0362 0.3653 $\pm$ 0.0694
fid $\downarrow$ 157.11 $\pm$ 49.91 183.24 $\pm$ 50.73 279.13 $\pm$ 42.53 145.94 $\pm$ 50.85
kid $\downarrow$ 0.0930 $\pm$ 0.0457 0.1120 $\pm$ 0.0493 0.1836 $\pm$ 0.0560 0.0848 $\pm$ 0.0496

### v3

@@ E.1 Discretization and Numerical Integration
Given endpoints $\bm{x}_{A}$ and $\bm{x}_{B}$, we represent the geodesic path as a sequence of $n+1$ points:
@@
where $\gamma_{0}=\bm{x}_{A}$ and $\gamma_{n}=\bm{x}_{B}$ are fixed, and the interior points $\{\gamma_{1},\ldots,\gamma_{n-1}\}$ are optimized.
@@
    Compute the velocity vector: $\bm{v}_{i}=\gamma_{i+1}-\gamma_{i}$
@@
    Calculate the midpoint: $\bm{m}_{i}=\frac{1}{2}(\gamma_{i+1}+\gamma_{i})$
@@
    Evaluate the score function at the midpoint: $\bm{s}_{i}=\bm{s}(\bm{m}_{i})$
@@
$\mathcal{E}[\gamma]\approx\frac{1}{2}\sum_{i=0}^{n-1}\|\bm{v}_{i}\|^{2}_{g(\bm{m}_{i})}=\frac{1}{2}\sum_{i=0}^{n-1}\left[\|\bm{v}_{i}\|^{2}+\lambda(\bm{s}_{i}^{T}\bm{v}_{i})^{2}\right]$ (13)
@@
E.2 Riemannian Adam Optimization
@@
3:Initialize: Moment vectors $\bm{m}_{i}=\bm{0}$, $\bm{v}_{i}=\bm{0}$ for $i\in\{1,\ldots,n-1\}$
[…]
@@
8:  for $i=1$ to n1n-1 do
9:   $\bm{g}_{i}\leftarrow\nabla_{\gamma_{i}}\mathcal{E}_{\text{total}}$
[…]
11:   $\bm{m}_{i}\leftarrow\beta_{1}\bm{m}_{i}+(1-\beta_{1})\widetilde{\bm{g}}_{i}$
12:   $\bm{v}_{i}\leftarrow\beta_{2}\bm{v}_{i}+(1-\beta_{2})\widetilde{\bm{g}}_{i}^{2}$
13:   $\hat{\bm{m}}_{i}\leftarrow\bm{m}_{i}/(1-\beta_{1}^{t})$ $\triangleright$ Bias correction
14:   $\hat{\bm{v}}_{i}\leftarrow\bm{v}_{i}/(1-\beta_{2}^{t})$
15:   $\bm{\eta}_{i}\leftarrow\alpha\cdot\hat{\bm{m}}_{i}/(\sqrt{\hat{\bm{v}}_{i}}+\epsilon)$ $\triangleright$ Update direction
@@
17:   $\gamma_{i}\leftarrow\gamma_{i}-\bm{\eta}_{i}$ $\triangleright$ Update position
[…]
@@
Riemannian Gradient: The Euclidean gradient $\bm{g}$ is converted to a Riemannian gradient $\widetilde{\bm{g}}$ using:
$\widetilde{\bm{g}}=g(\bm{x})^{-1}\bm{g}$ (17)
For our metric tensor $g(\bm{x})=\mathbf{I}+\lambda\cdot\bm{s}(\bm{x})\bm{s}(\bm{x})^{T}$, we can efficiently compute this using the Sherman-Morrison formula:
$\widetilde{\bm{g}}=\bm{g}-\frac{\lambda\cdot(\bm{s}^{T}\bm{g})\cdot\bm{s}}{1+\lambda\cdot(\bm{s}^{T}\bm{s})}$ (18)
@@
$\texttt{ParallelTransport}(\bm{m},\bm{x}_{\text{old}},\bm{x}_{\text{new}},g)=\bm{m}-\frac{1}{2}\cdot\langle\bm{m},\bm{x}_{\text{new}}-\bm{x}_{\text{old}}\rangle_{g}\cdot(\bm{x}_{\text{new}}-\bm{x}_{\text{old}})$ (19)
@@
Appendix F Geodesic Computation Algorithm
@@
1:Input: Points $\bm{p}$, $\bm{q}$, diffusion timestep $t$, number of segments $n$, scale parameter $\lambda$
2:Output: Geodesic path $\gamma=\{\gamma_{0},\gamma_{1},\ldots,\gamma_{n}\}$ with $\gamma_{0}=\bm{p}$, $\gamma_{n}=\bm{q}$
@@
4:Sample noise $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$
[…]
@@
[…]
@@
[…]
@@
17:  for $i=0$ to n1n-1 do
18:   $\bm{v}_{i}\leftarrow\gamma_{i+1}-\gamma_{i}$ $\triangleright$ Segment velocity
19:   $\bm{m}_{i}\leftarrow\frac{1}{2}(\gamma_{i+1}+\gamma_{i})$ $\triangleright$ Segment midpoint
20:   $\mathcal{E}\leftarrow\mathcal{E}+\frac{1}{2}\|\bm{v}_{i}\|^{2}+\frac{\lambda}{2}(\bm{s}(\bm{m}_{i})^{T}\bm{v}_{i})^{2}$ $\triangleright$ Energy contribution
@@
23:  $\mathcal{E}_{\text{smooth}}\leftarrow\lambda_{\text{smooth}}\sum_{i=1}^{n-1}\|\gamma_{i+1}-2\gamma_{i}+\gamma_{i-1}\|^{2}$ $\triangleright$ Smoothness
24:  $\mathcal{E}_{\text{mono}}\leftarrow\lambda_{\text{mono}}\sum_{i=1}^{n-1}\text{ReLU}(\|\gamma_{i}-\gamma_{n}\|-\|\gamma_{i-1}-\gamma_{n}\|)$ $\triangleright$ Monotonicity
@@
35:   $\hat{\gamma}_{i}\leftarrow$ original clean point ($\bm{p}$ or $\bm{q}$)
@@
37:   $\hat{\gamma}_{i}\leftarrow$ Denoise($\gamma_{i}$, from $t$ to 0)
@@ Appendix G Interpolation
1:Input: Points $\bm{p}$, $\bm{q}$, number of interpolation points $n$, timestep $t$
2:Output: Interpolated sequence $\{\bm{y}_{0},\bm{y}_{1},\ldots,\bm{y}_{n}\}$
@@
4:Sample noise $\bm{\epsilon}\sim\mathcal{N}(0,\mathbf{I})$ $\triangleright$ Same noise for consistency
[…]
@@
16:   $\bm{y}_{i}\leftarrow\bm{p}$ $\triangleright$ Use original clean endpoint
@@
18:   $\bm{y}_{i}\leftarrow\bm{q}$ $\triangleright$ Use original clean endpoint
@@
20:   $\bm{y}_{i}\leftarrow$ DenoisingDiffusion($\tilde{\gamma}_{i}$, from $t$ to 0)
@@
23:return $\{\bm{y}_{0},\bm{y}_{1},\ldots,\bm{y}_{n}\}$
Appendix H Extrapolation
@@
1:Initialize $\bm{x}_{\text{current}}\leftarrow\bm{q}$
@@
6:  Compute score at current point: $\bm{s}\leftarrow\mathbf{s}(\bm{x}_{\text{current}})$
7:  Compute direction: $\bm{d}\leftarrow(1-\varepsilon)\cdot\bm{m}+\varepsilon\cdot\bm{s}$
8:  Normalize: $\bm{d}\leftarrow\bm{d}/\|\bm{d}\|\cdot\text{step\_size}$
9:  Update position: $\bm{x}_{\text{next}}\leftarrow\bm{x}_{\text{current}}+\bm{d}$
10:  Update momentum: $\bm{m}\leftarrow\beta\cdot\bm{m}+(1-\beta)\cdot\bm{d}$
11:  $\bm{x}_{\text{current}}\leftarrow\bm{x}_{\text{next}}$
12:  Add $\bm{x}_{\text{current}}$ to extrapolation path
@@
Appendix I Computational Complexity Benchmark
We acknowledge that our method is more computationally intensive than baselines (LERP, SLERP, NoiseDiffusion), which are essentially fixed (non-adaptive) computations. However, we believe the computational overhead is reasonable considering the quality improvements and more importantly the data-driven nature.
Scalability. Despite the added complexity, our method scales well to large images and datasets. As demonstrated with Stable Diffusion 2.1 on 512×512 images (4×64×64 latents), the metric calculation scales nicely with image (latents) dimensionality.
Here are results averaged over 100 instantiations:
Method 5 points 10 points 20 points 30 points 50 points
LERP 0.22 $\pm$ 0.03 0.21 $\pm$ 0.05 0.67 $\pm$ 0.42 0.67 $\pm$ 0.44 0.74 $\pm$ 0.35
NoiseDiffusion 0.99 $\pm$ 0.25 1.23 $\pm$ 0.33 2.64 $\pm$ 0.60 3.37 $\pm$ 0.73 4.77 $\pm$ 0.72
SLERP 0.99 $\pm$ 0.25 0.89 $\pm$ 0.24 1.04 $\pm$ 0.48 1.11 $\pm$ 0.41 1.08 $\pm$ 0.50
Ours 6.16 $\pm$ 0.58 20.7 $\pm$ 9.8 58.1 $\pm$ 10.4 82.8 $\pm$ 10.4 144 $\pm$ 13
Table 5: Computation time for Rotated MNIST (32×32 pixels) in $\times 10^{-3}$ seconds.
Method 5 points 10 points 20 points 30 points 50 points
LERP 0.63 $\pm$ 0.33 0.65 $\pm$ 0.31 0.57 $\pm$ 0.32 0.67 $\pm$ 0.38 0.72 $\pm$ 0.42
NoiseDiffusion 1.21 $\pm$ 0.62 1.98 $\pm$ 0.73 2.62 $\pm$ 0.63 3.90 $\pm$ 0.81 6.85 $\pm$ 13.52
SLERP 1.09 $\pm$ 0.52 1.67 $\pm$ 0.90 1.05 $\pm$ 0.60 0.98 $\pm$ 0.40 1.03 $\pm$ 0.51
Ours 14.5 $\pm$ 3.3 30.6 $\pm$ 6.6 58.7 $\pm$ 7.5 85.3 $\pm$ 10.6 146 $\pm$ 14
Table 6: Computation time for Stable Diffusion (4×64×64 latent) in $\times 10^{-3}$ seconds.
Optimization cost. The most intensive part is geodesic optimization, which scales linearly with the number of points sampled ($n$) due to the energy functional (Riemannian integral).
Context. While baselines are faster, they offer no guarantees of following the learned data manifold. Computing geodesics on implicitly defined manifolds inherently requires optimization—there’s no analytical shortcut for complex learned manifolds.
Appendix J Rotated MNIST
After setting up the dataset as in Section 4.2, we trained a DDPM (Ho et al., 2020) Model in PyTorch (version 2.7 and cuda 12.6) by exploiting the Diffusers library (von Platen et al., 2022) and Accellerate library (Gugger et al., 2022) to parallelize training on 7 NVIDIA A100 GPUs.
@@ Appendix I Computational Complexity Benchmark
  • Input/Output: The model accepts $\times32$ grayscale images (single channel) and predicts noise with matching dimensions.
@@
Appendix K Stable Diffusion & MorphBench
We employed the pre-trained Stable Diffusion 2.1. implementation available in the Diffusers library (von Platen et al., 2022), running with PyTorch 2.7 and cuda 12.6 on a NVIDIA A100 GPU (40GB VRAM).
@@
For our experiments with Stable Diffusion 2.1, we used unconditional generation by providing an empty text prompt (””), allowing the model to focus solely on the image manifold structure without text-based guidance.
@@
Additionally we report more examples of interpolation comparison, see Figure 6,  7,  8
@@
avg_ssim $\uparrow$ 0.6251 $\pm$ 0.0108 0.6000 $\pm$ 0.0120 0.4202 $\pm$ 0.0115 0.6261 $\pm$ 0.0103
avg_psnr $\uparrow$ 21.23 $\pm$ 0.21 20.72 $\pm$ 0.24 17.09 $\pm$ 0.23 20.98 $\pm$ 0.21
avg_lpips $\downarrow$ 0.3467 $\pm$ 0.0075 0.3652 $\pm$ 0.0082 0.4577 $\pm$ 0.0058 0.3402 $\pm$ 0.0068
fid $\downarrow$ 125.89 $\pm$ 4.4 142.08 $\pm$ 4.9 253.52 $\pm$ 5.4 127.29 $\pm$ 5.0
kid $\downarrow$ 0.0947 $\pm$ 0.0065 0.1059 $\pm$ 0.0072 0.1925 $\pm$ 0.0079 0.0901 $\pm$ 0.0066
@@
avg_ssim $\uparrow$ 0.6282 $\pm$ 0.0129 0.5653 $\pm$ 0.0152 0.3969 $\pm$ 0.0120 0.6148 $\pm$ 0.0122
avg_psnr $\uparrow$ 21.29 $\pm$ 0.17 20.21 $\pm$ 0.22 16.58 $\pm$ 0.19 20.83 $\pm$ 0.17
avg_lpips $\downarrow$ 0.3672 $\pm$ 0.0077 0.4076 $\pm$ 0.0078 0.4891 $\pm$ 0.0036 0.3653 $\pm$ 0.0069
fid $\downarrow$ 157.11 $\pm$ 5.0 183.24 $\pm$ 5.1 279.13 $\pm$ 4.3 145.94 $\pm$ 5.1
kid $\downarrow$ 0.0930 $\pm$ 0.0046 0.1120 $\pm$ 0.0049 0.1836 $\pm$ 0.0056 0.0848 $\pm$ 0.0050